HETAV PATEL

Grade 11,

IF X-3Y=0 IS THE BISECTOR OF ONE ANGLE BETWEEN THE LINES 2x 2 -hxy+y 2 =0 then h=?

IF X-3Y=0 IS THE BISECTOR OF ONE ANGLE BETWEEN THE LINES 2x2-hxy+y2=0 then h=?

Grade:12

1 Answers

mohit sengar
13 years ago

the eq of bisector of the pair of homogeneous eq of lines is  x² - y²/ a-b = xy / h  if the eq is in the form ax² + 2hxy +by² = 0 now on putting values we get the eq of bisectors as   hx² +2xy -hy² =0     -(1)

we hav one eq as x - 3y =0 and let another eq as ax+by = 0 . 

on multiplying them we get the eq ax² + (b-3a)xy -3y² = 0    (2)

on comparing eq 1 nd 2 we get the value of h as  -3/4